OpenStudy (anonymous):

simplify the expression 3to the 4th power times 3 to the 3rd power

OpenStudy (unklerhaukus):

\[x^n\times x^m=x^{n+m}\] \[3^4\times3^3=3^{4+3}=3^{?}=\cdots\]

Parth (parthkohli):

I can explain how you get this property. Explanation is simple.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

21?

Parth (parthkohli):

\(3^3 \cdot 3^4 = 3 \cdot 3 \cdot 3\cdot 3 \cdot 3 \cdot 3 \cdot 3 = 3^?\)

Parth (parthkohli):

That is why you add the powers :)

OpenStudy (unklerhaukus):

\[x\times x=x^2\]\[x\times x\times x=x^3\]\[x\times x\times x\times x\times x=x^5=x^{2+3}\]

OpenStudy (anonymous):

it said the answer was 3 to the 7 or somethin like that but i keep getting 2187 or somthin like that i forgot

OpenStudy (unklerhaukus):

good work \[3^4\times3^3=3^7=2187\]

OpenStudy (unklerhaukus):

\[(3\times3\times3\times3)\times(3\times3\times3)=2187\]

OpenStudy (anonymous):

it says thats wrong thoee...its like three to the something power